The vertices of a triangle are P(-3,2), Q(-1,6) and R(4,1). Write an equation for the line containing each side.
Answer by stanbon(75887)

The vertices of a triangle are P(-3,2), Q(-1,6) and R(4,1). Write an equation for the line containing each side.
------------------
Line PQ:
slope = (6-2)/(-1--3) = 4/2 = 2
intercept: 2 = 2*-3 + b
b = 8
Equation: y = 2x + 8
==============================
Line QR:
slope = (1-6)/(4--1) = -5/5 = -1
intercept: 1 = -1*4 + b
b = 5
Equation: y = -x+5
=======================
I'll leave equation of PR to you.
Cheers,
Stan H.